# Heavy Fermions as an Efficient Representation of Atomistic Strain and Relaxation in Twisted Bilayer Graphene

### v1

@@ A.2 Graphene with Strain
and the orbital sites to be $\mathbf{r}_{1}=(\mathbf{a}_{1}-\mathbf{a}_{G,2})/3,\mathbf{r}_{2}=C_{6}\mathbf{r}_{1}$. The $K_{G}$-point is located at $\mathbf{K}=\frac{4\pi}{3a}(1,0)$. The nearest neighbor tight-binding model without strain is
@@ A.3 BM Model with Heterostrain
$\mathbf{\tilde{a}}^{l}_{G,i}$ $=\big{(}R(-l\theta/2)+\mathcal{E}_{l}\big{)}\mathbf{a}_{G,i}+O(\epsilon\theta)\ .$ (61)
@@
Rotating the strain axis by an angle $\varphi$ is accomplished by taking $\mathcal{E}^{l}\to R^{-1}(\varphi)\mathcal{E}^{l}R(\varphi)$, from which we find that $(\epsilon_{-},\epsilon_{xy})^{T}\to R(-2\varphi)(\epsilon_{-},\epsilon_{xy})^{T}$ transforms as a vector. For simplicity, we take $\varphi=0$ unless otherwise stated. Secondly, we also only consider heterostrains (opposite strain in each layer) such that $\mathcal{E}^{-}=-\mathcal{E}^{+}\equiv\mathcal{E}/2$. Homostrain (same strain in each layer) can be neglected to leading order as we will show momentarily.
@@
$\mathbf{\tilde{b}}_{G,i}^{l}$ $=\big{(}(R(-l\theta/2)+\mathcal{E}_{l})^{T}\big{)}^{-1}\mathbf{b}_{G,i}=\big{(}R(-l\theta/2)-\mathcal{E}^{T}_{l}\big{)}\mathbf{b}_{G,i}+O(\epsilon\theta)$ (64)
@@
$\mathbf{\tilde{K}}^{l}$ $=\big{(}R(-l\theta/2)-\mathcal{E}_{l}^{T}\big{)}\mathbf{K}_{G},\quad\mathbf{K}_{G}=\frac{4\pi}{3a}(1,0),\quad a=0.246\text{nm}$ (65)
@@
$\mathbf{\tilde{q}}_{j}$ $=\big{(}R(\theta/2)-R(-\theta/2)\big{)}\mathbf{K}_{j}-\mathcal{E}^{T}\mathbf{K}_{j},\qquad\mathbf{K}_{j}=C^{j-1}_{3}\mathbf{K}_{G}$ (66)
@@
$h_{K,l}(\mathbf{k})$ $=v_{F}(R(-l\theta/2)+\mathcal{E}^{l})^{-1}(\mathbf{k}-\mathbf{A}^{l})\cdot\boldsymbol{\sigma},\qquad\mathbf{A}^{l}=\frac{\beta}{a/\sqrt{3}}(\epsilon_{-}^{l},-\epsilon_{xy}^{l})$ (68)
@@
$\tilde{h}^{K}_{\mathbf{\tilde{Q}}\alpha,\mathbf{\tilde{Q}}^{\prime}\alpha^{\prime}}(\mathbf{k})=\delta_{\mathbf{\tilde{Q}},\mathbf{\tilde{Q}}^{\prime}}(v_{F}(\mathbf{k}-\mathbf{\tilde{Q}})+\frac{\beta}{a/\sqrt{3}}\zeta_{\mathbf{\tilde{Q}}}(\epsilon_{-}/2,\epsilon_{xy}/2))\cdot\boldsymbol{\sigma}_{\alpha\alpha^{\prime}}+\left(\sum_{i=1}^{3}\delta_{\mathbf{\tilde{Q}},\mathbf{\tilde{Q}}^{\prime}+\mathbf{\tilde{q}}_{j}}[T_{j}]_{\alpha\alpha^{\prime}}+h.c.\right)\ .$ (69)
@@ B.1 Relaxation and Strain
$D[C_{3}]h(\mathcal{E},\mathbf{k})D[C_{3}]^{\dagger}=h(C_{3}^{-1}\mathcal{E}C_{3},C_{3}\mathbf{k})$ (77)
since the strain tensor transforms as $\mathcal{E}_{l}\to R^{-1}(2\pi/3)\mathcal{E}_{l}R(2\pi/3)$ as shown after Eq. (63). This is equivalent to $(\epsilon_{-},\epsilon_{xy})^{T}\to R(-2\varphi)(\epsilon_{-},\epsilon_{xy})^{T}$ and $\epsilon_{+}\to\epsilon_{+}$. For example, we consider the $\Gamma_{3}$ block. By $C_{2z}\mathcal{T}$ and $C_{3z}$, we can write the allowed perturbations in terms of the unknowns $A(\mathcal{E}),B(\mathcal{E})$ as
$h_{\Gamma_{3}}(\mathcal{E},\mathbf{k}=0)=(A,B)\cdot(\sigma_{1},-\sigma_{2})$ (78)
where $A,B$ are first order in strain. The $C_{3}$ representation on this block is $D_{\Gamma_{3}}[C_{3}]=e^{\frac{2\pi i}{3}\sigma_{3}}$ which takes
$D_{\Gamma_{3}}[C_{3}]\begin{pmatrix}\sigma_{1}\\ -\sigma_{2}\end{pmatrix}D^{\dagger}_{\Gamma_{3}}[C_{3}]=R(-\frac{2\pi}{3})\begin{pmatrix}\sigma_{1}\\ -\sigma_{2}\end{pmatrix}=R(2\times\frac{2\pi}{3})\begin{pmatrix}\sigma_{1}\\ -\sigma_{2}\end{pmatrix}$ (79)
and thus $(A,B)$ must transform like $R^{T}(2\times\frac{2\pi}{3})=R(-2\times\frac{2\pi}{3})$. The only compatible components of the strain are $(A,B)\propto(\epsilon_{-},\epsilon_{xy})$. Repeating this procedure for the various blocks, we find the following parameterization of the perturbation $\delta H(\mathcal{E})$ to be
$\delta h(\mathcal{E})=\begin{pmatrix}c(\epsilon_{xy}\sigma_{1}-\epsilon_{-}\sigma_{2})&c^{\prime}(\epsilon_{xy}\sigma_{1}+\epsilon_{-}\sigma_{2})&i\gamma^{\prime}\epsilon_{+}\sigma_{3}\\ c^{\prime}(\epsilon_{xy}\sigma_{1}+\epsilon_{-}\sigma_{2})&M^{\prime}\epsilon_{+}\sigma_{2}&c^{\prime\prime}(\epsilon_{xy}\sigma_{0}+i\epsilon_{-}\sigma_{3})\\ -i\gamma^{\prime}\epsilon_{+}\sigma_{3}&c^{\prime\prime}(\epsilon_{xy}\sigma_{0}-i\epsilon_{-}\sigma_{3})&M_{f}(\epsilon_{xy}\sigma_{1}-\epsilon_{-}\sigma_{2})\\ \end{pmatrix}\ .$ (80)
@@
The parameters for strain and relaxation are calculated in the $\Gamma$ point approximation like the original HF parameters. Recall that the HF Hamiltonian $\delta h$ is computed from $\delta h_{\mu\nu}(\mathbf{k})=\braket{\mathbf{k},\mu}{H(\mathbf{k})}{\mathbf{k},\nu}$ where $\mu,\nu$ index the $f$- and $c$-mode wavefunctions and $H(\mathbf{k})$ is the BM or KV model in the plane wave basis. By construct, they are smooth and slowly varying around $\Gamma$. To compute any mass term (i.e. a $\mathbf{k}$-independent term), we simply have $\delta h(\mathbf{k}=0)=\braket{\mathbf{0},\mu}{H(\mathbf{0})}{\mathbf{0},\nu}$. Velocity terms are computed via $\boldsymbol{\nabla}_{k}\delta h(\mathbf{k})|_{\mathbf{k}=0}=\braket{\mathbf{0},\mu}{\boldsymbol{\nabla}_{k}\delta h(\mathbf{k})H(\mathbf{0})}{\mathbf{0},\nu}$ which is an approximation since it neglects the wavefunction derivatives $\boldsymbol{\nabla}_{\mathbf{k}}\ket{\mathbf{k},\nu}$. However, the smoothness of the wavefunctions ensures these derivatives are not large, and the approximation is confirmed numerically from the accuracy of the HF bands.

### v3

@@ A.2 Graphene with Strain
and the orbital sites to be $\mathbf{r}_{1}=(\mathbf{a}_{G,1}-\mathbf{a}_{G,2})/3,\mathbf{r}_{2}=C_{6}\mathbf{r}_{1}$. The $K_{G}$-point is located at $\mathbf{K}=\frac{4\pi}{3a}(1,0)$. The nearest neighbor tight-binding model without strain is
@@ A.3 BM Model with Heterostrain
$\mathbf{\tilde{a}}^{l}_{G,i}$ $=\big{(}R(l\theta/2)+\mathcal{E}_{l}\big{)}\mathbf{a}_{G,i}+O(\epsilon\theta)\ .$ (60)
@@
Rotating the strain axis by an angle $\varphi$ is accomplished by taking $\mathcal{E}^{l}\to R^{-1}(\varphi)\mathcal{E}^{l}R(\varphi)$, from which we find that $(\epsilon_{-},\epsilon_{xy})^{T}\to R(2\varphi)(\epsilon_{-},\epsilon_{xy})^{T}$ transforms as a vector. For simplicity, we take $\varphi=0$ unless otherwise stated. Secondly, we also only consider heterostrains (opposite strain in each layer) such that $\mathcal{E}^{-}=-\mathcal{E}^{+}\equiv\mathcal{E}/2$. Homostrain (same strain in each layer) can be neglected to leading order as we will show momentarily.
@@
$\mathbf{\tilde{b}}_{G,i}^{l}$ $=\big{(}(R(l\theta/2)+\mathcal{E}_{l})^{T}\big{)}^{-1}\mathbf{b}_{G,i}=\big{(}R(l\theta/2)-\mathcal{E}^{T}_{l}\big{)}\mathbf{b}_{G,i}+O(\epsilon\theta)$ (63)
@@
$\mathbf{\tilde{K}}^{l}$ $=\big{(}R(l\theta/2)-\mathcal{E}_{l}^{T}\big{)}\mathbf{K}_{G},\quad\mathbf{K}_{G}=\frac{4\pi}{3a}(1,0),\quad a=0.246\text{nm}$ (64)
@@
$\mathbf{\tilde{q}}_{j}$ $=\big{(}R(-\theta/2)-R(\theta/2)\big{)}\mathbf{K}_{j}-\mathcal{E}^{T}\mathbf{K}_{j},\qquad\mathbf{K}_{j}=C^{j-1}_{3}\mathbf{K}_{G}$ (65)
@@
$h_{K,l}(\mathbf{k})$ $=v_{F}(R(l\theta/2)+\mathcal{E}^{l})^{-1}(\mathbf{k}-\mathbf{A}^{l})\cdot\boldsymbol{\sigma},\qquad\mathbf{A}^{l}=\frac{\beta}{a/\sqrt{3}}(\epsilon_{-}^{l},-\epsilon_{xy}^{l})$ (67)
@@
$\tilde{h}^{K}_{\mathbf{\tilde{Q}}\alpha,\mathbf{\tilde{Q}}^{\prime}\alpha^{\prime}}(\mathbf{k})=\delta_{\mathbf{\tilde{Q}},\mathbf{\tilde{Q}}^{\prime}}(v_{F}(\mathbf{k}-\mathbf{\tilde{Q}})+\frac{\beta}{a/\sqrt{3}}\zeta_{\mathbf{\tilde{Q}}}(\epsilon_{-}/2,-\epsilon_{xy}/2))\cdot\boldsymbol{\sigma}_{\alpha\alpha^{\prime}}+\left(\sum_{i=1}^{3}\delta_{\mathbf{\tilde{Q}},\mathbf{\tilde{Q}}^{\prime}+\mathbf{\tilde{q}}_{j}}[T_{j}]_{\alpha\alpha^{\prime}}+h.c.\right)\ .$ (68)
@@ B.1 Relaxation and Strain
$D[C_{3}]h(\mathcal{E},\mathbf{k})D[C_{3}]^{\dagger}=h(C_{3}\mathcal{E}C_{3}^{-1},C_{3}\mathbf{k})$ (76)
since the strain tensor transforms as $\mathcal{E}_{l}\to R^{-1}(2\pi/3)\mathcal{E}_{l}R(2\pi/3)$ as shown after Eq. (62). This is equivalent to transforming $(\epsilon_{-},\epsilon_{xy})^{T}\to C_{3}(\epsilon_{-},\epsilon_{xy})^{T}$ and $\epsilon_{+}\to\epsilon_{+}$ in the Hamiltonian, placing constraints on the strain coupling in the heavy fermion basis. For example, we consider the $\Gamma_{3}$ block which transforms under the representation $D_{\Gamma_{3}}[C_{3}]=e^{\frac{2\pi i}{3}\sigma_{3}}$ which is a vector representation. Since $\boldsymbol{\sigma}=(\sigma_{1},\sigma_{2})$ also transforms in the vector representation of $D_{\Gamma_{3}}[C_{3}]$, the only symmetry-allowed combination of $\epsilon_{xy},\epsilon_{-}$ is $c\,\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}$. Next we consider the $\epsilon_{+}$ term in this block, which is invariant under $C_{3}$ (it transforms as a scalar). Both $\sigma_{0}$ and $\sigma_{3}$ transform as a scalar, but both commute with $D_{\Gamma_{3}}[P]=-i\sigma_{3}$, which must anti-commute with all strain terms since strain preserves particle-hole. Thus only the vector representation is allowed at leading order in $\mathbf{k}$.
Repeating this procedure for the various blocks, we find the following parameterization of the perturbation $\delta H(\mathcal{E})$ to be
$\delta h(\mathcal{E})=\begin{pmatrix}c(\epsilon_{xy}\sigma_{1}+\epsilon_{-}\sigma_{2})&c^{\prime}(\epsilon_{xy}\sigma_{1}-\epsilon_{-}\sigma_{2})&i\gamma^{\prime}\epsilon_{+}\sigma_{3}\\ c^{\prime}(\epsilon_{xy}\sigma_{1}-\epsilon_{-}\sigma_{2})&M^{\prime}\epsilon_{+}\sigma_{2}&c^{\prime\prime}(\epsilon_{xy}\sigma_{0}-i\epsilon_{-}\sigma_{3})\\ -i\gamma^{\prime}\epsilon_{+}\sigma_{3}&c^{\prime\prime}(\epsilon_{xy}\sigma_{0}+i\epsilon_{-}\sigma_{3})&M_{f}(\epsilon_{xy}\sigma_{1}+\epsilon_{-}\sigma_{2})\\ \end{pmatrix}\ .$ (77)
@@
The parameters for strain and relaxation are calculated in the $\Gamma$ point approximation like the original HF parameters. Recall that the HF Hamiltonian $\delta h$ is computed from $\delta h_{\mu\nu}(\mathbf{k})=\braket{\mathbf{k},\mu}{H(\mathbf{k})}{\mathbf{k},\nu}$ where $\mu,\nu$ index the $f$- and $c$-mode wavefunctions and $H(\mathbf{k})$ is the BM or KV model in the plane wave basis. By construction, they are smooth and slowly varying around $\Gamma$. To compute any mass term (i.e. a $\mathbf{k}$-independent term), we simply have $\delta h(\mathbf{k}=0)=\braket{\mathbf{0},\mu}{H(\mathbf{0})}{\mathbf{0},\nu}$. Velocity terms are computed via $\boldsymbol{\nabla}_{k}\delta h(\mathbf{k})|_{\mathbf{k}=0}=\braket{\mathbf{0},\mu}{\boldsymbol{\nabla}_{k}H(\mathbf{0})}{\mathbf{0},\nu}$ which is an approximation since it neglects the wavefunction derivatives $\boldsymbol{\nabla}_{\mathbf{k}}\ket{\mathbf{k},\nu}$. However, the smoothness of the wavefunctions ensures these derivatives are not large, and the approximation is confirmed numerically from the accuracy of the HF bands.